OpenStudy (anonymous):

5p+2=4p-1 solve for p and gve me the steps please:]

OpenStudy (anonymous):

5p - 4p = - 1 - 2 p = -3

OpenStudy (anonymous):

those rnt the steps m looking for...

OpenStudy (anonymous):

hhow did yu get -3?

OpenStudy (anonymous):

5p+2=4p-1 Since 4p contains the variable to solve for, move it to the left-hand side of the equation by subtracting 4p from both sides. 5p+2-4p=-1 Since 5p and -4p are like terms, add -4p to 5p to get p. p+2=-1 Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. p=-2-1 Subtract 1 from -2 to get -3. p=-3